OpenStudy (anonymous):

Find the solution to x = 3y – 8 and 4x + 5y = 19 by the substitution method. Would anyone mind giving me step by step instructions?

OpenStudy (anonymous):

x= 3y- 8- (1) 4x+ 5y= 19- (2) sub (1) into (2) 4(3y-8)+5y= 19 12y-32+5y= 19 17y-32= 19 17y= 51 y= 3 sub y=3 into (1) x= 3(3)-8= 1 therefore x=1, y=3

OpenStudy (anonymous):

Thank you so much! @jocelynong would you mind answering one more question for me?

OpenStudy (anonymous):

if i can, ill try

OpenStudy (anonymous):

@jocelynong Find the solution to 4x + y = 10 and 3x – y = 39 by the elimination method.

OpenStudy (anonymous):

it put question marks for some reason. it's 3x-y=39

OpenStudy (anonymous):

4x+y= 10- (1) 3x- y= 39 - (2) (2)+ (1): 3x-y+(4x+y)= 39+10 7x= 49 x= 7 so sub x= 7 in (1) 4(7)+ y= 10 y= -18 therefore x=7, y= -18

OpenStudy (anonymous):

thank youuuu! this means a lot to me :) @jocelynong

OpenStudy (anonymous):

hahah welcome. you don't happen to know differentiation do you? hehe